OpenStudy (anonymous):

Which of the following is the value of c for the best approximation for the quadratic function containing the points (5, 64), (7, 11), (9, -30), and (2, 4)? A. -73 B. 50 C. -5 D. -17

OpenStudy (mathmale):

There's some ambiguity in this question. Also, more points are given than are really necessary to fit a quadratic function as a best approximation. Given that we want a quadratic model, I'd immediately write out y=ax^2 + bx + c and then I'd assume that it is THIS c that we are supposed to find. Please brainstorm on what to do next. How would we find a, b and c? Perhaps focusing on c alone would save you some time.

OpenStudy (mathmale):

This problem is multi-step, by the way; you'll have to come up with 3 equations in 3 unknowns, namely, in a, b and c, and then you'll need to solve that system of linear equations, at least for c if not for a and b.

OpenStudy (ybarrap):

So, a quadratic equation looks like this f(x) =k(x - a)(x-b) We know that f(5)=64 and f(7) = 11 and f(2)=4 We have 3 equations and 3 unknowns. Solveable. There are the 3 equations: f(5)=64=k(5-a)(4-b)=k(20-5b-4a+ab) f(7)=11=k(11-a)(11-b)=k(120-11b-11a+ab) f(2)=4=k(2-a)(2-b)=k(4-2b-2a+ab) So we now have 20k-4kb-4ak+abk=64 120k-11kb-11ka+kab=11 3k-2kb-2ka+kab=4 Make sense?

OpenStudy (ybarrap):

I your case, you are looking for kab.

OpenStudy (mathmale):

Nice work, ybarrap. An alternative approach would be to use the model y=ax^2 + bx + c.

OpenStudy (anonymous):

@britneyspears